OpenStudy (rayman5766):

One hundred twenty is what percent of 200? A. 20% B. 60% C. 167% D. 240% I think the answer is A

OpenStudy (igreen):

120 = 200x Divide 200 to both sides.

OpenStudy (igreen):

What's 120 / 200?

OpenStudy (rayman5766):

0.6

OpenStudy (igreen):

Yes, now multiply by 100

OpenStudy (rayman5766):

60

OpenStudy (igreen):

Yep, that's our answer.

OpenStudy (rayman5766):

thanks I green I know I can all ways count on you

OpenStudy (mathstudent55):

200 is the whole and 120 is the part. You want to know what percent the part is of the whole. The whole is always 100%. You know this: whole is 100% and is 200 part is x% and is 120 Set up a proportion: \(\dfrac{100}{x} = \dfrac{200}{120} \) Now solve for x by cross multiplying: \(200x = 100 \times 200\) \(200x = 12,000\) \(x = \dfrac{12,000}{200} \) x = 60 Answer: 60%